# Eternal Universes

### v1

@@ Appendix D A Diffeomorphism of the Real Line
For the purpose of constructing a counter-example to the above conjecture one must be careful to preserve the required geodesic completeness of the scale factor. The scale factor must be defined for all $\mathbb{R}$ and have a strictly positive infimum. Thus, for the prescription to be a counter-example the diffeomorphism must compactify all $\mathbb{R}$ to the bounded interval where $-t^{2}+1>0$. Additionally, for the infinte to be compactified in a smooth fashion the derivative must approach zero on the asymptotes. The diffeomorphism, therefore, requires an inflection change. When said diffeomorphism is composed with the target concave down function, a balance of sort is created with the positive concavity of the diffeomorphism contrasting with the negative concavity of the target function. Although we have no formal proof this must always occur, in every calculation we have performed, we have always found at least a small interval where the positive inflection of the diffeomorphism wins out and the conjecture holds.
Finally, the smoothness assumption is important to rule out the piecewise linear case, where no concavity exists. Either a linear scale factor is unbroken and the positive definite requirement fails to hold or a broken linear scale factor fails to be smooth and contains singularities. A GFRW with a smooth scale factor is, by construction, free of curvature singularities. Thus, any non-trivial scale factor must be non-linear and can have at most a countable number of points with vanishing second derivative.

### v2

@@ D.3.2 Analytic Geodesics of the Quadratic Model
$x\left(t\right)$ $=\left(t-\frac{-a_{1}+\frac{\sqrt{a_{1}^{2}\omega_{0}-4a_{2}\sqrt{-\omega_{0}}-4a_{0}a_{2}\omega_{0}}}{\sqrt{\omega_{0}}}}{2a_{2}}\right)^{2}\times\cdots\times\left(\begin{matrix}\mathrm{complex\;solution\;of\;elliptic\;functions\;}\\ \mathrm{of\;variables\;}a_{2},a_{1},a_{0},\omega_{0},t\end{matrix}\ \right)$
$\qquad\;\times\cdots\times\left(\frac{-a_{1}+\frac{\sqrt{a_{1}^{2}\omega_{0}-4a_{2}\sqrt{-\omega_{0}}-4a_{0}a_{2}\omega_{0}}}{\sqrt{\omega_{0}}}}{2a_{2}}-\frac{-a_{1}+\sqrt{a_{1}^{2}+\frac{4a_{2}\left(\sqrt{-\omega_{0}}-a_{0}\omega_{0}\right)}{\omega_{0}}}}{2a_{2}}\right)+x_{0}$ (D.90)
with $\omega_{0}$ being the integration constant corresponding to causal character. Utilizing Eq. D.1 one reaps $\omega_{0}=\nicefrac{{1}}{{4}}$ corresponding to a timelike geodesic for initial conditions of $t_{0}=0$, $a_{2}=a_{1}=1$, $a_{0}=2$. Additionally, $x\left(t_{0}=0\right)=0$ demands $x_{0}\simeq 1.322$. Plot of this curve is shown in Figs. 26, 37, 38, 39.
Figure 36: Plot $x\left(t\right)$ vs $t$ of null and timelike geodesics of $\mathbb{R}^{1}_{1}\times_{a_{2}t^{2}+a_{1}t+a_{0}}\mathbb{R}^{3}$ with $a_{2}=a_{1}=1$ and $a_{0}=2$. Independent variable is the time coordinate. All length units have been renormalized to be dimensionless rescaling by canonical length scale $\sqrt{a_{2}}$. Initial null velocity is derived from frame $\left.\left[1,\frac{1}{2},0,0\right]^{T}\right|_{t=0}$ and initial timelike velocity is derived from frame $\left.\left[\sqrt{2},\frac{1}{2},0,0\right]^{T}\right|_{t=0}$ a propos Eq. D.1. The shown plot is a cylinder over constant coordinates $y,z$, hence two spacetime dimensions have been suppressed. Partial light cones can be seen as the area under the null curve shaded light gray. Small artifacts on plot are due to partial resolution of branch cuts in the complexified full analytic solution [^26] Full resolution of branch cuts of the analytic solution of Eq. D.3.2 is beyond the scope of this paper: this is left as an exercise to the reader. We would like to remind the reader, however, that Thm. 1 guarantees geodesic completeness of the solution, augmented by the complete numeric solutions of Figs. 33, 34, 35. See Tupper (2001) for more information.. Appropriate causal arclength was calculated to diverge as $t\rightarrow+\infty$; thus this model is geodesically complete. However, $H^{-}_{avg}=0$ for computation over first limit asymptotic past $t\rightarrow$-$\infty$, $H^{+}_{avg}=0$ for first limit asymptotic future $t\rightarrow+\infty$, and $H_{avg}>0$ for a symmetric compact interval limit computations - see Lesnefsky et al. (2023) for discussion of $H_{avg}$ limit order concerns. For the interval $\left[-10,10\right]$, one computes $H_{avg}^{\left[-10,10\right]}=\frac{\ln 28-\ln 23}{20}\simeq 0.009835$.
Figure 37: Additional future geodesic rays of the model and calculations of Fig. 26.
Figure 38: Additional past geodesic rays of the model and calculations of Fig. 26.
Figure 39: Additional geodesic plots of the model and calculations of Fig. 26. Additional null and timelike geodesics were plotted as to show the light cone cylinder $\Lambda_{\left(0,0,y,z\right)}\mathbb{R}^{1}_{1}\times_{a_{2}t^{2}+a_{1}t+a_{0}}\mathbb{R}^{3}$ over cylinder $t=0,x=0,y,z$.
Analytic null geodesics for the quadratic model can be calculated in a similar fashion to Sec. D.2.2. Geodesic null ODE can be calculated as
$\dot{x}\left(t\right)=\frac{1}{f\left(t\right)}=\frac{1}{a_{2}t^{2}+a_{1}t+a_{0}}$ (D.91)
with dimensionless solution
$\sqrt{a_{2}}x\left(\sqrt{a_{2}}t\right)=\frac{2\arctan\left(\frac{2\sqrt{a_{2}}t+\frac{a_{1}}{\sqrt{a_{2}}}}{\sqrt{4a_{0}-\frac{a_{1}^{2}}{a_{2}}}}\right)}{\sqrt{4a_{0}-\frac{a_{1}^{2}}{a_{2}}}}+\sqrt{a_{2}}\chi_{0}$ (D.92)
with integration constant $\chi_{0}$ corresponding to initial position $x_{0}$ vis a vis
$\chi_{0}=x_{0}-\frac{2\arctan\left(\frac{2\sqrt{a_{2}}t_{0}+\frac{a_{1}}{\sqrt{a_{2}}}}{\sqrt{4a_{0}-\frac{a_{1}^{2}}{a_{2}}}}\right)}{\sqrt{a_{2}}\sqrt{4a_{0}-\frac{a_{1}^{2}}{a_{2}}}}$ (D.93)
An analytic null geodesic solution for the quadratic model with parameters $a_{2}=a_{1}=1$ and $a_{0}=2$ is shown in Fig. 26.